For this exercise, give exact answers as simplified fractions. Compute E (X) and Var(X) if X has probability density function given by . . .

s c(1 – 26) if – 1< x <1 otherwise Determine the value of c as part of your answer.
Answer #1

1)for this to be valid:\int_{-1}^{1} f(x)dx =\int_{-1}^{1} c*(1-x6) dx =c*(x-x7/7) |1-1 =c*12/7=1

c=7/12

2)

E(X)=\int_{-1}^{1} xf(x)dx =\int_{-1}^{1} (7/12)*(x-x7) dx =c*(x2/2-x8/8) |1-1 =0

E(X2) =\int_{-1}^{1} x2f(x)dx =\int_{-1}^{1} (7/12)*(x2-x8) dx =c*(x3/3-x9/9) |1-1=7/27

therefore Var(X)=E(X2)-(E(X))2 =7/27-02=7/27